OpenStudy (momo):

What is the fifth term of a recursive sequence that follows the pattern 2x and has a first term x = 1? i think it is 2

myininaya (myininaya):

for x=1, 2x=2(1)=2 for x=2, 2x=2(2)=4 for x=3, 2x=2(3)=6 for x=4, 2x=2(4)=8 for x=5, 2x=2(5)=10 so you saying your pattern is 2,4,6,8,___ what goes in that blank, right? the 5th term is 10

myininaya (myininaya):

i think i don't really understand your question

myininaya (myininaya):

if your first term is 1, how could it follow a 2x pattern unless it started with x=1/2 but x=1 2x=1 when x=1/2 not x=1 so i'm not sure of the question

OpenStudy (momo):

the options are 2, 4, 8, and 16

OpenStudy (dumbcow):

i think 2x means multiply by 2 each time 1,2,4,8,16...

OpenStudy (dumbcow):

or an = 2^n

OpenStudy (momo):

well, to me, the answer to be 2 makes sense. i just want to know if i'm right lol

myininaya (myininaya):

well i don't think it would be 2 cow seems to have the right answer above

OpenStudy (dumbcow):

no the 5th term is 16 1,2,4,8,16...

OpenStudy (momo):

oh okay, i guess i misunderstood the question. thanks :)

OpenStudy (dumbcow):

if they are asking for common ratio then you are correct it is 2

OpenStudy (momo):

yeah, i think that's what i was confused with. thank you so much!